OpenStudy (anonymous):

find all real solutions 3x2 + 12x + 4 = 0

OpenStudy (nali):

factor then solve

OpenStudy (anonymous):

(3x+4)(x+1)=0?

OpenStudy (nali):

since u can not factor then just use the quardratic formula and canel the not real numbers

OpenStudy (nali):

http://www.regentsprep.org/Regents/math/algtrig/ATE3/quadformula.htm that is the quadratic formula

OpenStudy (nali):

use it and substutite a=3 b=12 and c=4

OpenStudy (nali):

did u get it?

OpenStudy (anonymous):

If you apply the quadratic formula, you will get \[ \left\{x=\frac{2}{3} \left(-3-\sqrt{6}\right),\quad x=\frac{2}{3} \left(\sqrt{6}-3\right)\right\} \]

OpenStudy (anonymous):

I didn't get it yet

OpenStudy (anonymous):

I got one = -.37 and the other as -3.63

OpenStudy (nali):

that is right